# Quasi-alternating surgeries

### v1

@@ 4. The asymmetric census knots
Among the SnapPy census L-space knots, there are exactly nine knots that are not strongly invertible. These are the asymmetric knots $\mathcal{A}$ studied in [ABG+23]:
@@ 4.1. Symmetry-exceptional slopes for the asymmetric census knots
Dunfield has already determined all non-hyperbolic filling slopes for knots in the SnapPy census [Dun20b]. We apply Strategy 3.3 to compute the symmetry-exceptional slopes of each knot in $\mathcal{A}$. ∎
@@ 4.2. Quasi-alternating slopes of the asymmetric census knots
Each knot in $\mathcal{A}$ admits exactly two dbc slopes, these two dbc slopes are consecutive integers, and each of these dbc slopes is quasi-alternating but not alternating. These slopes are presented in Tables 4.2 and 3.
Table 2. Information on the symmetric slopes of the asymmetric census L-space knots. A link marked with $(QA)$ is proven to be quasi-alternating.
@@
$(1,1)$ $\mathbb{Z}_{2}$ $-L(2,1)$ $L13n7360(0,0)(2,1)$
$(-1,1)$ $\mathbb{Z}_{2}$ $S^{3}$ $L12n789$ (QA)
$(-2,1)$ $\mathbb{Z}_{2}$ $-L(3,1)$ $L12n722(3,1)(0,0)$
$(-1,2)$ $\mathbb{Z}_{2}$ $L(3,1)$ $L11n192(-3,1)(0,0)$
$(-3,1)$ $\mathbb{Z}_{2}$ $-L(4,1)$ $L11n350(0,0)(-4,1)(0,0)$
@@
$(1,1)$ $\mathbb{Z}_{2}^{2}$bBigg@2.5{ $-L(5,1)$ $L11n419(0,0)(5,1)(0,0)$
$-L(3,1)$ $L12n1907(0,0)(3,1)(0,0)$
$-L(2,1)$ $L13n7356(0,0)(0,0)(2,1)$
$(-1,1)$ $\mathbb{Z}_{2}$ $S^{3}$ $L11n172$ (QA)
$(-1,2)$ $\mathbb{Z}_{2}$ $L(3,1)$ $L10a81(-3,1)(0,0)$
$(-1,3)$ $\mathbb{Z}_{2}$ $L(4,1)$ $L11a282(-4,1)(0,0)$
$(-2,3)$ $\mathbb{Z}_{2}$ $L(5,3)$ $L13n4363(-5,3)(0,0)$
$o9\_38928$
@@
$(-1,1)$ $\mathbb{Z}_{2}$ $L(3,1)$ $L12n703(-3,1)(0,0)$
$(-2,1)$ $\mathbb{Z}_{2}$ $L(2,1)$ $L12n1949(0,0)(2,1)(0,0)$
$(2,1)$ $\mathbb{Z}_{2}^{2}$bBigg@2.5{ $L(3,1)$ $L13n7433(0,0)(-3,1)(0,0)$
$L(2,1)$ $L10a105(0,0)(-2,1)$
$L(8,1)$ $L10n54(-8,1)(0,0)$
$o9\_39162$
@@
$(-1,1)$ $\mathbb{Z}_{2}$ $L(3,1)$ $L13n9864(-3,1)(0,0)(0,0)$
$(2,1)$ $\mathbb{Z}_{2}$ $-L(7,1)$ $L10n44(7,1)(0,0)$
$(1,2)$ $\mathbb{Z}_{2}$ $-L(3,1)$ $L12n784(3,1)(0,0)$
$(1,4)$ $\mathbb{Z}_{2}$ $L(7,4)$ $L13n4343(-7,4)(0,0)$
$o9\_40363$
@@
$(1,1)$ $\mathbb{Z}_{2}^{2}$bBigg@2.5{ $-L(2,1)$ $L14n43377(0,0)(2,1)$
$-L(3,1)$ $L13n9451(0,0)(3,1)(0,0)$
$-L(7,1)$ $L11n419(0,0)(7,1)(0,0)$
$(-1,1)$ $\mathbb{Z}_{2}$ $S^{3}$ $L12n785$ (QA)
$(-1,2)$ $\mathbb{Z}_{2}$ $L(3,1)$ $L11a225(-3,1)(0,0)$
$(-1,3)$ $\mathbb{Z}_{2}$ $-L(4,1)$ $L12a1861(0,0)(4,1)(0,0)$
$(-3,4)$ $\mathbb{Z}_{2}$ $L(7,4)$ $L13n4363(-7,4)(0,0)$
Table 3. Table 4.2 continued.
@@
$(-1,1)$ $\mathbb{Z}_{2}$ $L(3,1)$ $L12n715(-3,1)(0,0)$
$(-2,1)$ $\mathbb{Z}_{2}$ $L(2,1)$ $L12n1037(-2,1)(0,0)$
$(2,1)$ $\mathbb{Z}_{2}^{2}$bBigg@2.5{ $L(3,1)$ $L13n7421(0,0)(-3,1)(0,0)$
$-L(5,1)$ $L13n8521(0,0)(5,1)(0,0)$
$L(2,1)$ $L10a146(0,0)(-2,1)(0,0)$
$o9\_40582$
@@
$(-1,1)$ $\mathbb{Z}_{2}$ $L(5,3)$ $L10n44(-5,3)(0,0)$
$(1,2)$ $\mathbb{Z}_{2}$ $-L(3,1)$ $L12n996(3,1)(0,0)$
$(2,1)$ $\mathbb{Z}_{2}$ $-L(3,1)$ $L13n9833(3,1)(0,0)(0,0)$
$(4,1)$ $\mathbb{Z}_{2}$ $L(5,1)$ $L11n350(0,0)(-5,1)(0,0)$
$o9\_42675$
@@
$(1,1)$ $\mathbb{Z}_{2}$ $-L(5,3)$ $L14n24428(5,3)(0,0)$
$(-1,1)$ $\mathbb{Z}_{2}$ $S^{3}$ $K12n730$ (QA)
$(2,1)$ $\mathbb{Z}_{2}$ $L(2,1)$ $L13n7552(0,0)(-2,1)(0,0)$
$(-2,1)$ $\mathbb{Z}_{2}^{2}$bBigg@2.5{ $L(2,1)$ $L11n223(-2,1)(0,0)$
$-L(6,1)$ $L10n24(0,0)(6,1)$
$-L(8,3)$ $L13n9832(0,0)(8,3)(0,0)$
@@
In the first two cases, we perform the branch locus search (Strategy 3.5) to search for an explicit branching set if the symmetry group is $\mathbb{Z}_{2}$ and for three explicit branching sets if the symmetry group is $\mathbb{Z}_{2}^{2}$. In Tables 4.2 and 3 we present the simplest surgery description of the branching set that we found via this search. The symmetry exceptional slope is a dbc slope if and only if the quotient manifold for one of these symmetries is $S^{3}$.
@@
Finally, we show that no knot in $\mathcal{A}$ admits an alternating surgery. This fact was established by other techniques in [BKM24a]. We can immediately deduce that none of the symmetry-exceptional slopes are alternating using the uniqueness of the branching sets given by Lemma 2.3. Since each of the links in Tables 4.2 and 3 are non-alternating as evident from the n in their DT name, it only remains to argue that none of the three small Seifert fibered slopes is alternating. In general, small Seifert fibered spaces may arise as the double branched cover several different links in $S^{3}$. However, Motegi [Mot17, Theorem 3.1] applies to show that these specific Seifert fibered spaces arise only as the double branched covers of a unique link. ∎
@@ 5.1. The classification of the census L-space knots
Amongst the knot exteriors in the SnapPy census, there are exactly two knots where the L-space status is not decided [Dun20a]. These are the knots $o9\_30150$ and $o9\_31440$. Here we confirm that they are L-space knots by showing that they admit quasi-alternating surgeries.
@@
The knots $o9\_30150$ and $o9\_31440$ both have a quasi-alternating surgery. Hence they are L-space knots.
@@
We identify a quasi-alternating slope using Strategy 3.13. For that, we first use Strategy 3.5 to demonstrate that the $(-1,1)$-fillings on both of these manifolds are double branched coverings of small crossing knots:
@@
Figure 1(Left) shows a diagram of the knot $K13n3009$ with a crossing circled. This knot has determinant $87$. One smoothing of the knot gives the knot $K12n666$ which is quasi-alternating by Jablan’s table [Jab14] and has determinant $81$. The other smoothing gives the connected sum of a trefoil and a Hopf link which is alternating and has determinant $6$. Since $87=81+6$ and both smoothings are quasi-alternating, the knot $K13n3009$ is quasi-alternating.
Figure 1(Right) shows a diagram of the knot $K13n028$ with a crossing circled. This knot has determinant $105$. One smoothing of the knot gives the knot $K12n598$ which is quasi-alternating by Jablan’s table [Jab14] and has determinant $99$. The other smoothing gives the connected sum of a trefoil and a Hopf link which is alternating and has determinant $6$. Since $105=99+6$ and both smoothings are quasi-alternating, the knot $K13n2028$ is quasi-alternating. ∎
Figure 1. (Left) A diagram of the knot $K13n3009$ with a quasi-alternating crossing circled. (Right) A diagram of the knot $K13n2028$ with a quasi-alternating crossing circled.
@@ 5.2. Existence of quasi-alternating surgeries
Every census L-space knot admits a quasi-alternating surgery, except for $t09847$ and $o9\_30634$.
@@
  • Every census knot except for $t09847$ and $o9\_30634$ admits a quasi-alternating surgery. A quasi-alternating surgery for each census L-space knot without an alternating surgery or a quasi-alternating Seifert fibered space surgery is displayed in Table 4. (The alternating surgeries can be found in [BKM24a].)
Table 4. We display a quasi-alternating surgery for every census L-space knot that does not have an alternating surgery or a quasi-alternating Seifert fibered space surgery. The $17$ crossing knot $000815117nh\_$ 0008151 is in Burton’s notation [Bur20].
@@
For the first two bullets we load from Dunfield’s list of exceptional surgeries [Dun20b] all Seifert fibered space surgeries and use Issa’s classification of quasi-alternating Montesinos links [Iss18] to verify the statements.
@@
The remaining two knots are $t09847$ and $o9\_30634$ for which this search did not find a quasi-alternating branching set. In fact, these two knots admit no Khovanov thin surgery and hence no quasi-alternating surgery as shown in [BKM24b]. ∎
Table 5. The $8$ SFS surgeries along census L-space knots where the corresponding Montesinos links are not quasi-alternating.
@@ 6. The Baker–Luecke asymmetric L-space knots
In [BL20] an infinite family of interesting hyperbolic asymmetric L-space knots was constructed and in [BL20, BKM24a] their alternating slopes were studied. The subfamily of knots $K_{(m,b1,a1,a2,a3)}$ defined in [BL20, §11.4] for non-negative integers $a_{3},a_{2},a_{1},m,b_{1}$ gives a small collection of examples. In [BKM24a] we have shown that the $14$ simplest of these knots all have exactly two alternating slopes, and these pairs of slopes are consecutive integers. Here we show that $6$ of these have no further dbc slopes and give evidence that the same holds true for the other $8$ knots.
@@
    For the remaining $8$ of the above $14$ knots there are still two dbc surgeries. Both these dbc surgeries are integers and alternating. These surgeries together with their branching sets are displayed in Table 6. Moreover, for each of these $8$ knots there are at most two more slopes that might be dbc slopes. These further possible dbc slopes are present in Table 6. All other slopes are proven to not be dbc slopes.
Table 6. The first $14$ asymmetric L-space knots from [BL20], their symmetry-exceptional slopes, and their branching sets and quotient manifolds (if known). We observe that every knot has exactly two dbc slopes that are both alternating and consecutive integers. The slopes that are marked with unknown branching set are just conjectured (but not proven) to be not dbc slopes.
@@
$(417,1)$ $\mathbb{Z}_{2}$ $-L(3,1)$ $L14n16219(3,1)(0,0)$
$(831,2)$ $\mathbb{Z}_{2}$ $L(3,1)$ $L14a9663(-3,1)(0,0)$
(1,1,2,1,0)](1,1,2,1,0)]( 1 , 1 , 2 , 1 , 0 ) ] $(556,1)$ $\mathbb{Z}_{2}$ $S^{3}$ $L14a13430$
@@
$(1156,1)$ $\mathbb{Z}_{2}$ $S^{3}$ 15 crossing alternating link [BL20]
@@
$(1010,1)$ $\mathbb{Z}_{2}$ $S^{3}$ 15 crossing alternating link [BL20]
@@
$(966,1)$ $\mathbb{Z}_{2}$ $S^{3}$ 15 crossing alternating link [BL20]
@@
$(846,1)$ $\mathbb{Z}_{2}$ $S^{3}$ 15 crossing alternating link [BL20]
@@
$(844,1)$ $\mathbb{Z}_{2}$ $S^{3}$ 15 crossing alternating link [BL20]
@@
$(912,1)$ $\mathbb{Z}_{2}$ $S^{3}$ 15 crossing alternating link [BL20]
$(1,1,0,1,2)$ $(1142,1)$ $\mathbb{Z}_{2}$ $S^{3}$ 15 crossing alternating link [BKM24a]
@@
$(2,1,0,1,1)$ $(1066,1)$ $\mathbb{Z}_{2}$ $S^{3}$ 15 crossing alternating link [BKM24a]
@@
As in [BKM24a], we load the surgery description of the $K_{(m,b1,a1,a2,a3)}$ given in [BL20] into SnapPy from which we can build the $14$ examples shown in Table 6. Then the verified functions in SnapPy tell us that these knots are hyperbolic with trivial symmetry group. That they are L-space knots was shown in [BL20] by demonstrating that they have alternating surgeries.
@@
Then we use the branch locus search (Strategy 3.5) to look for the branching sets of the symmetry-exceptional slopes. Every one of these knots has exactly two alternating slopes; these were already found in [BKM24a]. Moreover, we have identified explicit branching sets for $5$ more slopes among these knots. These other $5$ branching sets are given by links in manifolds different from $S^{3}$ and thus do not certify these slopes as dbc slopes. Since $4$ of these $5$ slopes have symmetry $\mathbb{Z}_{2}$, this is sufficient to conclude that those slopes are not dbc slopes. The remaining slope is the one with symmetry $\mathbb{Z}_{2}^{2}$ and its other branching sets have not yet been determined.
For the other symmetry-exceptional slopes our code could not identify explicit branching sets. This is most likely because the branching sets are too complicated to appear in the list over which we have iterated. Nevertheless, we expect these slopes to not be dbc slopes. But at the moment we cannot prove that and thus they appear as possibly dbc slopes in Table 6. ∎

### v2

@@ 4. The asymmetric census knots
Among the SnapPy census L-space knots, there are exactly nine knots that are not strongly invertible. These are the asymmetric knots $\mathcal{A}$ studied in [ABG+19]:
@@
Recall, that for the census manifolds we display slopes in the geometric basis as used by SnapPy. From the homology of the fillings it is straightforward to express the geometric basis in terms of the Seifert basis. This is displayed in Table 1.
Table 1. For each asymmetric census L-space knot we display the relation between the slopes $r_{S}$ measured with respect to the Seifert basis $(\mu_{S},\lambda_{S})$ and the slopes $r_{G}$ measured with respect to the geometric basis $(\mu_{G},\lambda_{G})$. For all census L-space knots it turns out that $\mu_{S}=\mu_{G}$.
knot slope knot slope knot slope
$t12533$ $r_{S}=r_{G}+37$ $t12681$ $r_{S}=r_{G}+61$ o9_38928o9\_38928 $r_{S}=r_{G}+50$
o9_39162o9\_39162 $r_{S}=r_{G}+65$ o9_40363o9\_40363 $r_{S}=r_{G}-83$ o9_40487o9\_40487 $r_{S}=r_{G}+38$
o9_40504o9\_40504 $r_{S}=r_{G}-58$ o9_40582o9\_40582 $r_{S}=r_{G}+44$ o9_42675o9\_42675 $r_{S}=r_{G}+46$
@@ 4.1. Symmetry-exceptional slopes for the asymmetric census knots
Dunfield has already determined all non-hyperbolic filling slopes for knots in the SnapPy census [Du18]. We apply Strategy 3.3 to compute the symmetry-exceptional slopes of each knot in $\mathcal{A}$. ∎
@@ 4.2. Quasi-alternating slopes of the asymmetric census knots
Each knot in $\mathcal{A}$ admits exactly two dbc slopes, these two dbc slopes are consecutive integers, and each of these dbc slopes is quasi-alternating but not alternating. These slopes are presented in Tables 3 and 4.
Table 3. Information on the symmetric slopes of the asymmetric census L-space knots. A link marked with $(QA)$ is proven to be quasi-alternating. If a symmetry group is marked with $SFS$ the surgered manifold is proven to be a Seifert fibered space.
@@
$(1,1)$ $\mathbb{Z}_{2}$ L(2,1)-L(2,1) $L13n7360(0,0)(2,1)$
(1,1)(-1,1) $\mathbb{Z}_{2}$ $S^{3}$ $L12n789$ (QA)
(2,1)(-2,1) $\mathbb{Z}_{2}$ L(3,1)-L(3,1) $L12n722(3,1)(0,0)$
(1,2)(-1,2) $\mathbb{Z}_{2}$ $L(3,1)$ L11n192(3,1)(0,0)L11n192(-3,1)(0,0)
(3,1)(-3,1) $\mathbb{Z}_{2}$ L(4,1)-L(4,1) L11n350(0,0)(4,1)(0,0)L11n350(0,0)(-4,1)(0,0)
@@
[…]
L(3,1)-L(3,1) $L12n1907(0,0)(3,1)(0,0)$
L(2,1)-L(2,1) $L13n7356(0,0)(0,0)(2,1)$
(1,1)(-1,1) $\mathbb{Z}_{2}$ $S^{3}$ $L11n172$ (QA)
(1,2)(-1,2) $\mathbb{Z}_{2}$ $L(3,1)$ L10a81(3,1)(0,0)L10a81(-3,1)(0,0)
(1,3)(-1,3) $\mathbb{Z}_{2}$ $L(4,1)$ L11a282(4,1)(0,0)L11a282(-4,1)(0,0)
(2,3)(-2,3) $\mathbb{Z}_{2}$ $L(5,3)$ L13n4363(5,3)(0,0)L13n4363(-5,3)(0,0)
o9_38928o9\_38928
@@
(1,1)(-1,1) $\mathbb{Z}_{2}$ $L(3,1)$ L12n703(3,1)(0,0)L12n703(-3,1)(0,0)
(2,1)(-2,1) $\mathbb{Z}_{2}$ $L(2,1)$ $L12n1949(0,0)(2,1)(0,0)$
[…]
$L(2,1)$ L10a105(0,0)(2,1)L10a105(0,0)(-2,1)
$L(8,1)$ L10n54(8,1)(0,0)L10n54(-8,1)(0,0)
o9_39162o9\_39162
@@
(1,1)(-1,1) $\mathbb{Z}_{2}$ $L(3,1)$ L13n9864(3,1)(0,0)(0,0)L13n9864(-3,1)$(0,0)$
$(2,1)$ $\mathbb{Z}_{2}$ L(7,1)-L(7,1) $L10n44(7,1)(0,0)$
$(1,2)$ $\mathbb{Z}_{2}$ L(3,1)-L(3,1) $L12n784(3,1)(0,0)$
$(1,4)$ $\mathbb{Z}_{2}$ $L(7,4)$ L13n4343(7,4)(0,0)L13n4343(-7,4)(0,0)
o9_40363o9\_40363
@@
[…]
L(3,1)-L(3,1) $L13n9451(0,0)(3,1)(0,0)$
L(7,1)-L(7,1) $L11n419(0,0)(7,1)(0,0)$
(1,1)(-1,1) $\mathbb{Z}_{2}$ $S^{3}$ $L12n785$ (QA)
(1,2)(-1,2) $\mathbb{Z}_{2}$ $L(3,1)$ L11a225(3,1)(0,0)L11a225(-3,1)(0,0)
(1,3)(-1,3) $\mathbb{Z}_{2}$ L(4,1)-L(4,1) $L12a1861(0,0)(4,1)(0,0)$
(3,4)(-3,4) $\mathbb{Z}_{2}$ $L(7,4)$ L13n4363(7,4)(0,0)L13n4363(-7,4)(0,0)
Table 4. Table 3 continued.
@@
(1,1)(-1,1) $\mathbb{Z}_{2}$ $L(3,1)$ L12n715(3,1)(0,0)L12n715(-3,1)(0,0)
(2,1)(-2,1) $\mathbb{Z}_{2}$ $L(2,1)$ L12n1037(2,1)(0,0)L12n1037(-2,1)(0,0)
[…]
L(5,1)-L(5,1) $L13n8521(0,0)(5,1)(0,0)$
$L(2,1)$ L10a146(0,0)(2,1)(0,0)L10a146(0,0)(-2,1)(0,0)
o9_40582o9\_40582
@@
(1,1)(-1,1) $\mathbb{Z}_{2}$ $L(5,3)$ L10n44(5,3)(0,0)L10n44(-5,3)(0,0)
$(1,2)$ $\mathbb{Z}_{2}$ L(3,1)-L(3,1) $L12n996(3,1)(0,0)$
$(2,1)$ $\mathbb{Z}_{2}$ L(3,1)-L(3,1) $L13n9833(3,1)(0,0)(0,0)$
$(4,1)$ $\mathbb{Z}_{2}$ $L(5,1)$ L11n350(0,0)(5,1)(0,0)L11n350(0,0)(-5,1)(0,0)
o9_42675o9\_42675
@@
$(1,1)$ $\mathbb{Z}_{2}$ L(5,3)-L(5,3) $L14n24428(5,3)(0,0)$
(1,1)(-1,1) $\mathbb{Z}_{2}$ $S^{3}$ $K12n730$ (QA)
$(2,1)$ $\mathbb{Z}_{2}$ $L(2,1)$ L13n7552(0,0)(2,1)(0,0)L13n7552(0,0)(-2,1)(0,0)
[…]
L(6,1)-L(6,1) $L10n24(0,0)(6,1)$
L(8,3)-L(8,3) $L13n9832(0,0)(8,3)(0,0)$
@@
In the first two cases, we perform the branch set search (Strategy 3.5) to search for an explicit branching set if the symmetry group is $\mathbb{Z}_{2}$ and for three explicit branching sets if the symmetry group is $\mathbb{Z}_{2}^{2}$. In Tables 3 and 4 we present the simplest surgery description of the branching set that we found via this search. The symmetry exceptional slope is a dbc slope if and only if the quotient manifold for one of these symmetries is $S^{3}$.
@@
Finally, we show that no knot in $\mathcal{A}$ admits an alternating surgery. This fact was established by other techniques in [BKM_alt]. We can immediately deduce that none of the symmetry-exceptional slopes are alternating using the uniqueness of the branching sets given by Lemma 2.4. Since each of the links in Tables 3 and 4 are non-alternating as evident from the n in their DT name, it only remains to argue that none of the three small Seifert fibered slopes is alternating. In general, small Seifert fibered spaces may arise as the double branched cover of several different links in $S^{3}$. However, Motegi [Mo17, Theorem 3.1] applies to show that these specific Seifert fibered spaces arise only as the double branched covers of a unique link, which is the Montesinos link associated to the Seifert invariants of the manifold. ∎
@@ 5.1. The classification of the census L-space knots
Amongst the knot exteriors in the SnapPy census, there are exactly two knots where the L-space status is not decided [Du19]. These are the knots o9_30150o9\_30150 and o9_31440o9\_31440. Here we confirm that they are L-space knots by showing that they admit quasi-alternating surgeries.
@@
The knots o9_30150o9\_30150 and o9_31440o9\_31440 both have a quasi-alternating surgery. Hence they are L-space knots.
@@
We identify a quasi-alternating slope using Strategy 3.13. For that, we first use Strategy 3.5 to demonstrate that the (1,1)(-1,1)-fillings on both of these manifolds are double branched coverings of small crossing knots:
@@
Figure 1(Left) shows a diagram of the knot $K13n3009$ with a crossing circled. This knot has determinant $87$. One smoothing of the knot gives the knot $K12n666$ which is quasi-alternating by Jablan’s table [Ja14] and has determinant $81$. The other smoothing gives the connected sum of a trefoil and a Hopf link which is alternating and has determinant $6$. Since $87=81+6$ and both smoothings are quasi-alternating, the knot $K13n3009$ is quasi-alternating.
Figure 1(Right) shows a diagram of the knot $K13n2028$ with a crossing circled. This knot has determinant $105$. One smoothing of the knot gives the knot $K12n598$ which is quasi-alternating by Jablan’s table [Ja14] and has determinant $99$. The other smoothing gives the connected sum of a trefoil and a Hopf link which is alternating and has determinant $6$. Since $105=99+6$ and both smoothings are quasi-alternating, the knot $K13n2028$ is quasi-alternating. ∎
Figure 1. (Left) A diagram of the knot $K13n3009$ with a quasi-alternating crossing circled. (Right) A diagram of the knot $K13n2028$ with a quasi-alternating crossing circled. Both diagrams were created using SnapPy [CDGW].
@@ 5.2. Existence of quasi-alternating surgeries
Every census L-space knot admits a quasi-alternating surgery, except for $t09847$ and o9_30634o9\_30634.
@@
  • Every census knot except for $t09847$ and o9_30634o9\_30634 admits a quasi-alternating surgery. A quasi-alternating surgery for each census L-space knot without an alternating surgery or a quasi-alternating Seifert fibered space surgery is displayed in Table 5. (The alternating surgeries can be found in [BKM_alt].)
Table 5. We display a quasi-alternating surgery for every census L-space knot that does not have an alternating surgery or a quasi-alternating Seifert fibered space surgery. The $17$ crossing knot 17nh_000815117nh\_0008151 is in Burton’s notation [Burton].
@@
For the first two bullets we load from Dunfield’s list of exceptional surgeries [Du18] all Seifert fibered space surgeries and use Issa’s classification of quasi-alternating Montesinos links [Is18] to verify the statements.
@@
The remaining two knots are $t09847$ and o9_30634o9\_30634 for which this search did not find a quasi-alternating branching set. In fact, these two knots admit no Khovanov thin slope and hence no quasi-alternating slope as shown in [BKM_thin]. ∎
Table 6. The 88 SFS surgeries along census L-space knots where the corresponding Montesinos links are not quasi-alternating. Here we are using the Regina notation for Seifert fibered spaces [BBP+]. Refer to Section 7 for more details.
@@ 6. The Baker–Luecke asymmetric L-space knots
In [BL17] an infinite family of interesting hyperbolic asymmetric L-space knots was constructed and in [BL17, BKM_alt] their alternating slopes were studied. The subfamily of knots $K_{(m,b1,a1,a2,a3)}$ defined in [BL17, §11.4] for non-negative integers $a_{3},a_{2},a_{1},m,b_{1}$ gives a small collection of examples. In [BKM_alt] we have shown that the $14$ simplest of these knots all have exactly two alternating slopes, and these pairs of slopes are consecutive integers. Here we show that $6$ of these have no further dbc slopes and give evidence that the same holds true for the other $8$ knots.
@@
    For the remaining $8$ of the above $14$ knots there are still two dbc surgeries. Both these dbc surgeries are consecutive integers and alternating. These surgeries together with their branching sets are displayed in Table 7. Moreover, for each of these $8$ knots there are at most two more slopes that might be dbc slopes. These further possible dbc slopes are present in Table 7. All other slopes are proven to not be dbc slopes.
Table 7. The first $14$ asymmetric L-space knots from [BL17], their symmetry-exceptional slopes, and their branching sets and quotient manifolds (if known). We observe that every knot has exactly two dbc slopes that are both alternating and consecutive integers. The slopes that are marked with unknown branching set are just conjectured (but not proven) to be not dbc slopes.
@@
$(417,1)$ $\mathbb{Z}_{2}$ L(3,1)-L(3,1) $L14n16219(3,1)(0,0)$
$(831,2)$ $\mathbb{Z}_{2}$ $L(3,1)$ L14a9663(3,1)(0,0)L14a9663(-3,1)(0,0)
$(1,1,2,1,0)]$ $(556,1)$ $\mathbb{Z}_{2}$ $S^{3}$ $L14a13430$
@@
$(1156,1)$ $\mathbb{Z}_{2}$ $S^{3}$ 15-crossing alternating link [BL17]
@@
$(1010,1)$ $\mathbb{Z}_{2}$ $S^{3}$ 15-crossing alternating link [BL17]
@@
$(966,1)$ $\mathbb{Z}_{2}$ $S^{3}$ 15-crossing alternating link [BL17]
@@
$(846,1)$ $\mathbb{Z}_{2}$ $S^{3}$ 15-crossing alternating link [BL17]
@@
$(844,1)$ $\mathbb{Z}_{2}$ $S^{3}$ 15-crossing alternating link [BL17]
@@
$(912,1)$ $\mathbb{Z}_{2}$ $S^{3}$ 15-crossing alternating link [BL17]
$(1,1,0,1,2)$ $(1142,1)$ $\mathbb{Z}_{2}$ $S^{3}$ 15-crossing alternating link [BKM_alt]
@@
$(2,1,0,1,1)$ $(1066,1)$ $\mathbb{Z}_{2}$ $S^{3}$ 15-crossing alternating link [BKM_alt]
@@
As in [BKM_alt], we load the surgery description of the $K_{(m,b1,a1,a2,a3)}$ given in [BL17] into SnapPy from which we can build the $14$ examples shown in Table 7. Then the verified functions in SnapPy tell us that these knots are hyperbolic with trivial symmetry group. That they are L-space knots was shown in [BL17] by demonstrating that they have alternating surgeries.
@@
Then we use the branch set search (Strategy 3.5) to look for the branching sets of the symmetry-exceptional slopes. Every one of these knots has exactly two alternating slopes; these were already found in [BKM_alt]. Moreover, we have identified explicit branching sets for $5$ more slopes among these knots. These other $5$ branching sets are given by links in manifolds different from $S^{3}$ and thus do not certify these slopes as dbc slopes. Since $4$ of these $5$ slopes have symmetry $\mathbb{Z}_{2}$, this is sufficient to conclude that those slopes are not dbc slopes. The remaining slope is the one with symmetry $\mathbb{Z}_{2}^{2}$ and its other branching sets have not yet been determined. For the other symmetry-exceptional slopes our code could not identify explicit branching sets. These are marked as ‘unknown’ in the table. ∎
Remark 6.2
The fact that the branch set search did not terminate for some of these examples is because the branching sets are too complicated to appear in our search. We expect that these slopes are not dbc slopes, that is, we expect the quotient manifold is not $S^{3}$. However as the search was inconclusive, their status as dbc slopes remains unknown in Table 7.